A students is going to graduate from a business department of a university by the end of the semester. After being interviewed at two companies he likes, he assesses that his probability of getting an offer from company A is 0.6, and his probability of getting an offer from company B is 0.45. If he believes that the probability that he will get offers from both companies is 0.35, what is the probability that he will get at least one offer from these two companies?
1
Expert's answer
2020-06-12T10:21:19-0400
"P(A\cup B)=P(A)+P(B)-P(A\cap B)"

"P(A\cup B)=0.6+0.45-0.35=0.7"

The probability that he will get at least one offer from these two companies is 0.7